OpenStudy (anonymous):

Which does not represent y as a function of x? x = y^2 + 2 y = x^2 + 2 x = y + 8 y = -x + 8 Need help :/

OpenStudy (aum):

x = y^2 + 2 The same x value such as x = 3 can produce two different y values: When x = 3, 3 = y^2 + 2 y^2 = 1 y = -1, +1. The same x gives two different y values. It fails the vertical line text and therefore it is not a function.